OpenStudy (anonymous):

The cost to produce a custom designed t-shirt is five dollars per shirt plus a 40 dollar design fee. What would be the cost to produce 50 shirts?

OpenStudy (anonymous):

you always have to pay 40 dollar design. then every shirt costs 5 dollar. so for 50 shirts you must pay 50*5$+40$=250$+40$=290$

OpenStudy (anonymous):

Equation for shirts (x represents t shirts): 5x+40 50 tshirts sub into x =50*5x+40 =290 dollars

OpenStudy (anonymous):

why 50*5x[sic]+40 ? i think you subbed the x? ;)